### v1

# CNN on ‘Top’: In Search of Scalable & Lightweight Image-based Jet Taggers

Abstract
While Transformer-based and standard Graph Neural Networks (GNNs) have proven to be the best performers in classifying different types of jets, they require substantial computational power. We explore the scope of using a lightweight and scalable version of EfficientNet architecture, along with global features of the jet. The end product is computationally inexpensive but is capable of competitive performance. We showcase the efficacy of our network in tagging top-quark jets in a sea of other light quark and gluon jets.
journal: Eur. Phys. J. C
1 Introduction
As we enter the high-energy (HE) and high luminosity (HL) phase of the Large Hadron Collider (LHC), the study of jet physics has become more relevant than ever. ‘Fat jet tagging’ is a critical task for identifying the hadronic decays of massive particles, such as the Higgs boson, top quark, or a $W$ or $Z$-boson that are produced with high momentum or ‘boost’ in a particular direction. As a result of this boost, the decay products of these standard model (SM) particles are highly collimated and can be contained within a single large-radius jet, which we call a ‘fat jet’. Scenarios like this can easily occur in various beyond the standard model (BSM) frameworks with vector-like quarks Yang:2025ktj ; Ghosh:2025gdq , leptoquarks Sahoo:2025kdj ; Bhaskar:2024snl ; Ghosh:2025gue , heavy right-handed neutrinos Chakraborty:2018khw ; Padhan:2022fak , supersymmetric particles with $R$-parity violating couplings Baruah:2024wrn ; Sahu:2025uop , to name a few. Among the above-mentioned SM particles, top quark holds a unique status. It happens to be the heaviest fundamental particle we know thus far, and unlike other quarks, it decays before hadronization. The heavy mass of the top quark means its coupling strength with the Higgs boson is the strongest. Hence, precise measurements of top quark properties provide a consistency check of the SM. Even subtle deviations from the SM predictions for top quark interactions could hint at new physics. Moreover, the top quark mass plays a crucial role in determining the stability of the electroweak vacuum. Though the current measurements suggest that our vacuum is ‘metastable’ Hiller:2024zjp ; Degrassi:2012ry ; Buttazzo:2013uya , this conclusion is highly sensitive to the precise values of the top quark measurements Alekhin:2012py ; Elias-Miro:2011sqh ; Espinosa:2015kwx . Identifying top quark jets as fat jets in a sea of other light quark and gluon jets, therefore, is a rather important task.
Machine learning (ML) has revolutionized this field Sahu:2024fzi ; Bhattacherjee:2022gjq ; Mikuni:2025tar by moving beyond traditional methods that relied on high-level, physics-inspired features like jet mass, jet radius, $n$-subjettiness, energy correlation functions, etc Thaler:2010tr ; Thaler:2011gf ; Stewart:2010tn ; Larkoski:2013eya ; Moult:2016cvt ; Komiske:2017aww ; Abdesselam:2010pt ; Kogler:2018hem ; Altheimer:2012mn ; Altheimer:2013yza ; Adams:2015hiv ; Marzani:2019hun ; Larkoski:2017jix . Deep learning models, on the other hand, instead of using simple, hand-crafted features, train themselves to represent the jets as points in a vector space of features best suited for the subsequent classification. The inputs to these models include the four-momentum components of each constituent particle, global features that carry the kinematical properties of the jet, and particle tracking IDs whenever available. From the ML perspective, the jet physics task at hand can be divided into two broad categories: classification and generation. Both of these ‘downstream’ tasks depend on representing the raw, low-level particle or detector data associated with a jet as points in a dense, continuous vector space (feature space), ready to be effectively processed by any ML algorithm. This process, known as ‘embedding/encoding,’ allows the model to learn a more comprehensive and nuanced representation of the substructure of a jet, which can be utilized in subsequent downstream tasks.
[…]
A more direct and fundamental approach to embedding is to represent the jet as a point cloud of its constituent particles. Each ‘point’ is a particle with its own set of features, e.g., four momentum components, particle IDs, charges, and tracking information. GNNs Shlomi:2020gdn ; Duarte:2020ngm ; Thais:2022iok are particularly well-suited for this representation. The jet is now treated as a graph where each particle is a ‘node’, and the connections between them are ‘edges’ Komiske:2018cqr ; Qu:2019gqs ; Bogatskiy:2020tje ; Gong:2022lye ; Bogatskiy:2023nnw ; Bogatskiy:2022czk ; Bogatskiy:2023fug . Constituent information can also be represented as a varying length sequence, especially if it is ordered based on a certain property, e.g., by their transverse momentum $p_{T}$. RNNs—originally developed for processing text or speech—can then be used. The advent of the self- and cross-attention mechanisms gave us the famously parallelizable and scalable Transformer architecture DBLP:journals/corr/VaswaniSPUJGKP17 . Originally used to process and generate natural language sequences, these models can be thought of as special variations of GNNs JoshiTrans2GNN , and can process jets either as sequences of constituents or unordered point clouds.
One jarring issue that plagues the existing networks with the highest accuracy in identifying various jets is the computational cost. Achieving high accuracy boils down to extracting maximal information from the jet representation. It can either be gleaned by formidably large networks or can be preprocessed and fed to a smaller network later. The former is the path taken by transformer-based models Qu:2022mxj ; Mikuni:2021pou ; Wu:2024thh and more costly foundational models like Omnilearn(ed) Mikuni:2025tar ; Mikuni:2024qsr ; Bhimji:2025isp . Even these, when striving for higher accuracies, bring separately extracted features to the networks ‘attention’ [^1] Not adding pairwise information to the multi-head attention of the (Par-T) network results in a $1.2\%$ drop in accuracy and a $20-30\%$ drop in background rejection Qu:2019gqs ; Qu:2022mxj .. In the latter approach, when jets are considered as particle clouds or graphs, creating a fully connected graph is prohibitively costly in terms of time and computation, mainly because the pairwise interactions (edge information222i.e., the abstract message-passing, not physical QCD interactions.) are numerous for a jet with a large number of constituents. These challenges are either bypassed by keeping only the nearest neighbor interactions (ParticleNet Qu:2019gqs ), or by representing data in a geometric algebra over space-time, equivariant under physics-motivated (Lorentz) transformations Bogatskiy:2022czk ; Bogatskiy:2023nnw ; Gong:2022lye . The newer architectures combine both of these approaches Spinner:2024hjm ; Brehmer:2024yqw . Even running these state-of-the-art models on single-GPU machines becomes an ordeal, let alone training them from scratch. With the rising awareness of the virtues of Transfer Learning, more models are now pre-trained on larger datasets with multiple classes, which automatically requires significant model complexity. This considerably increases their performance, but the downside is that the cost of retraining increases many times over for specific downstream tasks.
Our objective in this work is to revisit computationally low-cost CNNs, as they implicitly utilize local, pairwise information (albeit spatially, through images) to some extent. Depending on the knowledge that the inclusion of global jet features enhances tagging efficiency Bhattacherjee:2022gjq , we explore the possibility of constructing a lightweight yet efficient network by incorporating global features with the images. Although computationally lightweight, smaller CNNs such as LeNet lenetPaper cannot achieve competitive accuracies. Extracting more information here is equivalent to increasing image resolution and/or model complexity (depth/width). Although CNNs become notoriously computationally costly when scaled up (e.g., ResNet ResNet1 ), recent developments, such as compound scaling EffNetV1 ; EffNetV2 , combined with depthwise and pointwise convolutionsMobileNetV1 ; MobileNetV2 ; MobileNetV3 ; MobileNetV4 ; MNASnet , resulted in networks that reach higher accuracies at a fraction of the cost. As all versions of EfficientNet were originally designed for images with higher resolution, we have built lightweight versions of those, while following the same scaling rules. We aim to explore their capabilities with and without the externally provided global features, while using LeNet as a benchmark for comparison throughout.
The paper is organized as follows. In Section 2 we provide a detailed account of our methodology. This includes machine configuration, training data preparation, model architectures, and training procedure. In Section 3 we present our results. Finally, in Section 4, we summarize our work and draw conclusions.
(a) $32\times 32$
(b) $32\times 32$, Standardized
(c) $32\xrightarrow{Crop}28$, Standardized
(d) $40\xrightarrow{Crop}32$, Standardized
(e) $32\times 32$
(f) $32\times 32$, Standardized
(g) $32\xrightarrow{Crop}28$, Standardized
(h) $40\xrightarrow{Crop}32$, Standardized
Figure 1: Representative images of two jets for different resolutions and pixel standardization. Although images in our dataset are rendered in RGB, we present them here using the CEI 1931 XYZ color scheme xyzcolorCEI:1931 , for better visual understanding. All images from the top panel are of one representative top jet, and those below are of one QCD jet. The first column shows the jet constituents as a $32\times 32$ image in the $\Delta\eta-\Delta\phi$ plane, as the channels are combined and rendered. The second panel in the same row contains the same image, with the pixels standardized (the mean image of the dataset is subtracted from the original and then divided by the standard deviation image). The third column displays the images in the second, but cropped to $28\times 28$. The fourth column contains the same jets, first rendered in $40\times 40$, then cropped to $32\times 32$.
Figure 2: Schematics of network architectures used in this work. On the left is LeNet, used as the benchmark CNN model. On the Right is the simplest Efficient Net (small) used. Representative schematics of the M-Conv blocks and the squeeze and expand residual blocks within those are also shown.
2 Methodology
2.1 Machine Configuration
All training and measurements, except for the measurement of inference speed, are performed on a single desktop PC with $64GB$ RAM, a $13^{th}$ generation Intel© i9 processor, and a $12GB$ NVIDIA© RTX A2000 graphics processor. The entire analysis is conducted in Mathematica® with the Wolfram Language, which uses the Apache MXNet® framework in the backend. The original EfficientNet architectures (before modification) were taken from the Wolfram Net Repository effnet:Wolfram . To compare the inference speed of CNNs in this work with some external models, inference times were calculated on another machine with an $11GB$ NVIDIA© GeForce GTX 1080-Ti graphics processor.
2.2 Data: Images from Constituents
We have used a pre-existing dataset [^3] https://syncandshare.desy.de/public.php/dav/files/llbX3zpLhazgPJ6/?accept=zip that contains both top-quark and mixed light-quark-gluon jets. The objective, as mentioned before, is to identify top quark jets as signals and tag them as fat jets amid the sea of other jets. The events were generated with 14 TeV center-of-mass energy using Pythia8 Sjostrand:2014zea followed by detector simulation performed using Delphes deFavereau:2013fsa . The jets were constructed using the anti-kT algorithm Cacciari:2008gp with FastJet Cacciari:2011ma . The pseudorapidity $|\eta_{j}|$ is required to be <2<2 for the fat top quark jet, whereas the radius parameter is chosen to be $0.8$. It is ensured that a matched parton-level top exists within $\Delta R=0.8$ of this signal jet. Moreover, all decay products of the top quark are required to be within $\Delta R=0.8$ of the jet axis. 4-momenta information is stored for the leading $200$ jet constituents. The leading top-quark jets lie within a $p_{T}$ range of 550650550-650 GeV. The total sample consists of one million top jets alongside another million non-top jets, which are treated here as background. The dataset is divided into training / validation / test ($1.2M/0.4M/0.4M$) sets.
We pre-process the data to create 3-channel images with transverse momentum ($p_{T}$), mass ($m$), and energy ($E$) in each channel, respectively. To create each channel from the jet constituents, the hardest constituent is considered to be the center of the channel, and the rest of the image is constructed by binning the remaining constituents in two dimensions: the difference in pseudo-rapidity ($\Delta\eta$) and that in the azimuthal angle ($\Delta\phi$). The paper associated with the original dataset Kasieczka:2019dbj assumes the calorimeter resolution to be $0.04\times 2.25^{\circ}$, where a fat jet with a radius parameter $R=0.8$ can be covered with $40\times 40$ pixels. Fixing the bin sizes thus keeps our jet images at this dimension or less ($32\times 32$). Due to these choices and the nature of the dataset, both $\Delta\eta$ and $\Delta\phi$ axes are retained from 1.2-1.2 to $1.2$, without the loss of any constituents (see Figs (1(a)) and (1(e)) ). Unlike the original paper, we do not fix the axis of the second hardest object; rather, we use an image augmentation layer to randomly flip the image along either axis or both during training. We do not use random cropping, as the jets are constructed objects and can always be centered around the hardest $p_{T}$ constituent for all practical applications. Furthermore, to use N-subjettiness later in our analysis, we also require each jet to have at least 4 constituents, which drops a meager 33 jets from the $201.8K$ dataset.
Standardization of input data has become common in ML, as it has been found to improve the performance of generic neural networks. We perform this standardization by creating a ‘mean’ and a ‘standard deviation’ image using a large number of random samples from the dataset. We respectively subtract from and divide every sample by the values of these arrays. This clearly isolates a fairly circular region in the $\Delta\eta-\Delta\phi$ plane containing the bulk of the jet structure (see Figs (1(b)) and (1(f)) ). During some preliminary runs, it has been consistently observed that using images where the central region covers most of the image field yields better classification performance. Thus, we apply a center-crop to the $32\times 32$ images, making them $28\times 28$ (see Figs (1(c)) and (1(g)) ). A similar size of the central region is obtained for the $40\times 40$ images when they are cropped to $32\times 32$ (see Fig.s (1(d)) and (1(h)) ). We keep more of the field of images than is usual practice in the community, where final images are generally cropped to $R=0.8$ Kasieczka:2017nvn ; Kasieczka:2019dbj ; Macaluso:2018tck . Our crops vary within $0.96\leq R\leq 1.05$.
2.3 Data: Jet-Level Information
We have combined some global features of the jet to ascertain how much they improve the tagging efficiency of the CNNs. The features used in this work are inspired by ref.s ATL-PHYS-PUB-2021-028 ; Bhattacherjee:2022gjq . The original dataset provides the four-momentum of up to 200 constituents per jet, from which all global features can be computed. For this purpose, we employ the fastjet package along with EnergyCorrelator contrib from fastjet-contrib-1.3.2.
In addition to the global four-momentum of the jet, our feature set includes the N-subjettiness variables Thaler:2010tr , as well as several series of observables constructed from Energy Correlation Functions Larkoski:2013eya and generalized correlator functions Moult:2016cvt . The complete list of variables is as follows:
  • Jet four-momentum: $p_{T}$, $\eta$, $\phi$, $m_{\text{jet}}$
  • Number of constituents: $N_{\text{cons}}$
  • NN-subjettiness ratios: $\tau_{21}^{\beta}$, $\tau_{32}^{\beta}$, $\tau_{43}^{\beta}$, $\tau_{54}^{\beta}$, $\tau_{65}^{\beta}$ for $\beta=0.5$, $1.0$, $2.0$
  • $C$-series: $C_{1}^{\beta}$, $C_{2}^{\beta}$, $C_{3}^{\beta}$ for $\beta=1.0,\,2.0$
  • $D$-series [^4] The parameter choices for the D-series features follow Larkoski:2014gra ; Larkoski:2015kga ; Larkoski:2014zma .: $D_{2}^{\alpha,\beta}$ for $\alpha,\beta=1.0,\,2.0$, and the four $D_{3}$ variables defined in Bhattacherjee:2022gjq
  • U-series: $U_{1}^{\beta},U_{2}^{\beta},U_{3}^{\beta}$ for $\beta=0.5,\,1.0,\,2.0$
  • M-series: $M_{2}^{\beta},M_{3}^{\beta}$ for $\beta=1.0,\,2.0$
  • N-series: $N_{2}^{\beta},N_{3}^{\beta}$ for $\beta=1.0,\,2.0$
  • L-series: $L_{2},L_{3}$ as defined in ATL-PHYS-PUB-2021-028
Certain global features, such as $U_{3}^{\beta}$ and $M_{3}^{\beta}$, require at least 4 constituents for their computation. As mentioned in the previous sub-section, we discard all jets with $N_{\text{constituents}}<4$. For complete definitions and a detailed description of these variables, we refer the reader to ref.s Bhattacherjee:2022gjq ; ATL-PHYS-PUB-2021-028
2.4 Model Architecture
Historically, one of the most successful convolutional neural networks (CNNs), LeNet lenetPaper , is also the oldest and one of the simplest. LeNet-5, with alternating convolutional and pooling layers, has been the dependable standard for smaller images (see the left side of Fig. (2) for its architecture). Over the decades following LeNet, CNNs continued to scale up to classify images at increasingly higher resolutions. AlexNet AlexNetPaper , Inception nets InceptionV1 ; InceptionV2V3 , ResNets ResNet1 and their wider versions wideResNet , DenseNets DenseNet , and MobileNets MobileNetV1 ; MobileNetV2 ; MobileNetV3 ; MobileNetV4 are only a few examples of these. To go deeper and wider, these newer architectures used varied techniques, such as ‘residual’ connections to feed information from earlier layers to the latter parts of the network, point-wise $(1\times 1)$ and depthwise convolutions for dimension reduction to remove computational bottlenecks, batch normalization, and mixtures of these methods InceptionV4 ; ResNext . Targeting general-purpose image classification, these models were optimized for the ImageNet dataset ImageNet2019 ; ImageNet2021 and its equivalents; thus, computationally, they are quite expensive.
Layers Input Resolution
$3\times 40\times 40$ $3\times 64\times 64$
Conv $3\times 3$ $16\times 20\times 20$ $16\times 32\times 32$
Block 1A $8\times 20\times 20$ $8\times 32\times 32$
Block 2A $8\times 10\times 10$ $16\times 16\times 16$
Block 3A $16\times 5\times 5$ $16\times 8\times 8$
Block 4A $24\times 3\times 3$ $32\times 4\times 4$
Block 5A $32\times 3\times 3$ $48\times 4\times 4$
Block 6A $56\times 2\times 2$ $80\times 2\times 2$
Block 7A $96\times 2\times 2$ $136\times 2\times 2$
Conv $3\times 3$ $400\times 2\times 2$ $536\times 2\times 2$
Aggregate $400$ $536$
FC & Out 22 22
Table 1: Effnet-S configurations for two different input resolutions. For more details, see sec. 2.4.
Network Input Image No. of Inference Time Accuracy AUC Bkg. Rej at TPR
Type Array Size Resolution Param.s /100 Jets ($ms$) (%) (%) $50\%$ $30\%$
LeNet $3\times 28\times 28$ $32\to 28$ $428K$ $5.5(7)$ $92.52$ $97.81$ $160(2)$ $525(23)$
(Images) $3\times 32\times 32$ $653K$ $7.3(8)$ $92.54$ $97.80$ $162(2)$ $536(14)$
$3\times 40\times 40$ $1253K$ $6.5(20)$ $92.56$ $97.82$ $169(3)$ $568(31)$
$3\times 32\times 32$ $40\to 32$ $653K$ $6.7(6)$ $92.55$ $97.82$ $165(4)$ $516(20)$
$3\times 40\times 40$ $1253K$ $6.5(7)$ $92.51$ $97.75$ $155(2)$ $476(16)$
$3\times 28\times 28$ $32\to 28$ $485K$ $6.4(6)$ $93.12$ $98.11$ $225(4)$ $774(33)$
$3\times 32\times 32$ $710K$ $7.5(5)$ $93.10$ $98.11$ $222(6)$ $777(35)$
(Images + $3\times 40\times 40$ $1310K$ $7.3(7)$ $93.08$ $98.11$ $222(3)$ $795(23)$
Global Feat.s) $3\times 32\times 32$ $40\to 32$ $710K$ $7.0(7)$ $93.17$ $98.14$ $233(2)$ $793(19)$
$3\times 40\times 40$ $1310K$ $7.7(8)$ $\mathbf{93.20}$ $\mathbf{98.15}$ $237(3)$ $\mathbf{814(24)}$
EffNet-S $3\times 40\times 40$ $32\to 28$ $162K$ $7.7(7)$ $92.59$ $97.85$ $174(5)$ $\mathbf{602(14)}$
(Images) $3\times 64\times 64$ $305K$ $10.4(6)$ $\mathbf{92.70}$ $\mathbf{97.91}$ $170(13)$ $577(67)$
$3\times 40\times 40$ $40\to 32$ $162K$ $8.5(6)$ $91.97$ $97.51$ $129(9)$ $415(40)$
(Images + $3\times 40\times 40$ $32\to 28$ $208K$ $8.4(6)$ $93.12$ $98.13$ $233(6)$ $\mathbf{864(44)}$
Global Feat.s) $3\times 64\times 64$ $365K$ $12.0(11)$ $\mathbf{93.13}$ $98.15$ $226(8)$ $817(38)$
DeepTop Macaluso:2018tck ; Kasieczka:2017nvn ; Kasieczka:2019dbj $1\times 37\times 37$ $60\to 40\to 37$ $610K$ - $93.00$ $98.10$ - $914(14)$
ResNeXt-50 ResNext
Ref. Kasieczka:2019dbj $1\times 64\times 64$ $96\to 64$ $1460K$ - $93.60$ $98.40$ - $1122(47)$
Ref. Qu:2019gqs $1\times 64\times 64$ $96\to 64$ $1460K$ 22 $93.60$ $98.37$ $302(5)$ $1147(58)$
Table 2: Single number performance metrics for ML models performed on the total test dataset. From the second to the seventh column, we respectively report the input array sizes, the resolution of the generated jet images ($A\to B$ means the original image was $A\times A$, and it was cropped to $B\times B$), total number of parameters in the network, inference time for 100 jets in milliseconds, overall accuracy, and area under the ROC curve (AUC). The last two columns list the QCD background rejection at top tagging efficiencies of $50\%$ and $30\%$ respectively.
Though scaling CNNs according to image resolution has been attempted before long2015fully , EfficientNets EffNetV1 ; EffNetV2 were the first to solve the scaling problem by using ‘compound scaling’, connecting network depth and width to image resolution. Given a baseline network (found from a neural architecture search, or NAS), these quantities scale as:
$\text{depth coefficient }d$ $=\alpha^{\phi}$
$\text{width coefficient }w$ $=\beta^{\phi}$
$\text{resulution coefficient }r$ $=\gamma^{\phi}\,$, (1)
where $\alpha\geq 1$, $\beta\geq 1$, $\gamma\geq 1$. Here, increasing $\phi$, the scaling parameter, increases the complexity of the network in a compounded manner. Noting that doubling the net-depth doubles the computational cost (floating point operations per second, or FLOPS), but doubling the other two quadruples the FLOPS, these also follow an empirical constraint: $\alpha\cdot\beta^{2}\cdot\gamma^{2}\simeq 2$, so that increasing $\phi$ by some $\phi_{0}$ increases the FLOPS by $2^{\phi_{0}}$. The authors of the first EfficientNet paper performed a grid search on $\alpha$, $\beta$, and $\gamma$ and obtained the values $1.2$, $1.1$, and $1.15$, respectively. Starting with $\phi=0$ and increasing $\phi$ by 1 each time, they obtained eight networks ($B_{0}$ to $B_{7}$) with increasing complexity. The base network ($B_{0}$) works on arrays of dimension $224\times 224$, while the largest of these models ($B_{7}$) takes in arrays of dimension $600\times 600$. There are seven main building blocks of an EfficientNet, each containing multiple parts. Building upon other successful lightweight models MobileNetV2 ; MNASnet , EfficientNets replace a fully convolutional operation with a depthwise convolution (a single filter per channel) and a pointwise ($1\times 1$) convolution (see the blown up schematic of Block 2A in Fig. (2) ). This significantly reduces the computational cost while maintaining similar performance. When stacked together, the blocks mentioned above also perform a function known as inverted bottlenecking, embedding the input manifold into a lower dimensional manifold. Each of these blocks also contains a squeeze and expansion block squeeze , with a residual connection maintaining the information flow, as well as a combination of Swish, RELU, and Sigmoid activation functions. More details can be found in the original paper EffNetV1 . While the width coefficient scales up the channel dimension of each block, the depth coefficient determines the number of times each type of block is repeated, making the network deeper. These networks also scale up the dropout probability to regularize the larger models. All versions of EfficientNet outperform other CNNs that are much larger in size (in terms of accuracy) while being faster during inference ($B_{0}$, with 5.3M5.3~M parameters, achieves a top-5 accuracy of $93.3\%$ for ImageNet, compared to the widely used ResNet-50 (26M26~M parameters)—which has a top-5 accuracy of $93.0\%$; the FLOPS for ResNet-50 is $11$ times that of EfficientNet-$B_{0}$).
We begin with LeNet in this analysis as a benchmark for performance. Following its historical usage, the smallest resolution of jet images we use is $28\times 28$, increasing to $32\times 32$ and $40\times 40$. Images larger than these make the jet constituents sparser, and CNNs with even higher-resolution inputs are needed for their analysis. The network we focus on here is EfficientNet. While they have been used on smaller images, such as the CIFAR-10 and CIFAR-100 datasets, these have been upsampled to fit the $224\times 224$ input dimensions  [^5] Author’s answer: https://github.com/tensorflow/tpu/issues/421.. This defeats one of the main purposes of this work (working with constrained computational resources), both in terms of RAM and VRAM.
Figure 3: ROC curve (signal efficiency vs. background rejection) for some of the networks (with some metric highest among others; table (2) has them in bold-face). The solid lines are for models working on both images and global features, dashed ones correspond to those with only images as input.
To circumvent this without changing the hyperparameters of the original EfficientNet architecture (and using the prescribed values of $\alpha$, $\beta$, and $\gamma$), we use negative $\phi$ values to scale down the input resolution. The smallest resolution thus achieved is $36\times 36$. Any smaller, and the bottlenecks within the network would not hold without modification. We choose the starting resolution to be $40\times 40$ (in accordance with the resolution of the simulated calorimeter), corresponding to $\phi\simeq-12.33$. A larger version was also created with $64\times 64$ inputs, for $\phi\simeq-8.96$ [^6] Similar efforts can be found here: https://github.com/qubvel/efficientnet/issues/26. The other important parameter to set is the dropout probability, which is gradually increased from $B_{0}$ to $B_{7}$, as more regularization is needed for larger networks. We linearly extrapolated these values using the width and depth coefficients and used the corresponding values for our networks. Evidently, the much smaller size of our networks resulted in significantly less dropout regularization. Altogether, we will denote this smaller architecture as EfficientNet-Small (EffNet-S) for this work. For these values of $\phi$, every block appears only once in the smaller EffNet-S structures. We have thus indexed the blocks ‘$i$-A’ ($i=1,\,2,\,\ldots$), to keep the possibility of larger networks open [^7] This raises an obvious question: as these CNNs were fine-tuned to detect specific shapes (with detail) from larger images, unlike in our case—where the jet-images are smaller and are mere representations with a much larger importance to finer structures—could it be possible to systematically search for and create versions of scalable networks for this specific task, albeit smaller? Though intriguing, this question is outside the scope of the present work. However, glimpses of this possibility may be perceived from the present analysis.. Table (1) lists the input dimensions for this architecture for two different input resolutions: $40\times 40$ and $64\times 64$.
2.5 Training
Although we load the validation and test datasets into RAM as a whole, due to our choice of confining ourselves to a single mid-range PC with $64GB$ RAM, we use a type of data piping for the training dataset. We load $300,000$ training data at a time, which we will call a ‘round’ from now on. Each round is trained until the validation loss stops decreasing for 10 epochs. An image of the trained network is saved from each round and used in the next one. This is done for 15 (18) rounds for $32\times 32$ ($40\times 40$) image representation of jets to ensure that the entire dataset is adequately represented. For most of the models used in this work, we repeat the process, starting with the last trained network, and use the same rounds again. This way, the rounds mimic the behavior of epochs, while storing the datasets beforehand reduces computation time. We then evaluate the accuracy of each of these models on the test dataset, and the one with the highest accuracy is chosen as the best version of that specific model architecture. All further measurements are performed on this iteration of the model.
To ensure optimal computation time, we choose batch sizes depending on whether the memory dedicated to the graphics card can handle them without overflowing into the SWAP memory. For EffNet-S (64) a batch size of $2048$ is used. For all other networks, it is fixed at $4096$. We use the ADAM optimizer kingma2014adam with default settings ($\beta_{1}=0.9$, $\beta_{2}=0.99$, and $\epsilon=10^{-5}$, along with the default learning rate update schedule) to minimize the cross-entropy loss, where we wait for a fixed number of epochs (generally 10) for the last minimum loss to be updated; otherwise, we stop the training.
The global features of the jets are standardized and utilized in the second part of the analysis, in conjunction with the images. These are concatenated with the output of the ‘Flatten Layer’ of LeNet and, in the case of EffNet-S, that of the ‘Aggregation Layer’, as can be seen from Fig. (2). Pre-trained networks are faster to retrain and reach higher accuracies after modification—a fact that has been re-established in recent attempts at foundational models in jet applications. Keeping that in mind, we use the first part of the best network of a particular architecture and add an MLP block to it in this part of the analysis.
We retrain the best network of each architecture nine additional times, with a representative subset of the dataset, to obtain the uncertainties arising from random weight initialization.
3 Results
Table (2) lists all single-number performance metrics for the networks explored in this analysis. The table is divided into three blocks: the first block contains all the information about the LeNet architecture, the second one about EffNet-S, and the third one lists some of the publicly available information about CNNs used previously as top-taggers for comparison. We have obtained results for LeNet-type networks with three different input resolutions: $28\times 28$, $32\times 32$, and $40\times 40$. Versions of EffNet-S are of two types: $40\times 40$ and $64\times 64$. Images are generated in two ways: $32\times 32$ images cropped to $28\times 28$ and $40\times 40$ images cropped to $32\times 32$, as explained in sec. 2.2. Each network is run on all images with cropped resolution (col. 3 of tab. 2) less than or equal to its input dimension (col. 2 of tab. 2). Networks with a specific architecture and input resolution optimized for maximum accuracy are used to create versions that can also process the global features.
LeNets with increasing input array-sizes, when applied only to images cropped to $28\times 28$, gain marginally in performance across all metrics. Incorporating global features clearly increases the discriminative power—especially background rejection—of the networks, and the effect is more prominent for higher image resolutions. This is expected, as LeNet is a conventional CNN. Full convolutions with a large kernel size ($5$) extract more information from the images, albeit at a higher computational cost and a heavier network. The LeNet-40 is $2.9(2.7)$ times the size of LeNet-40 without(with) global features.
The smallest EffNet-S, while applied only to images, outperforms all LeNets, while being $1/8$-th the size of the largest LeNet and in terms of the total number of parameters. Its performance is best when the input images are $32\xrightarrow{Crop}28$, and it drops a little for the larger images. The possible reason is that the ‘MB-Conv’ blocks with bottlenecks in EffNet-S do not perform at the level of full convolutions when analyzing sparse pixel-level information, and networks with larger input arrays are needed for proper classification. Adding global features is beneficial here as well, but the effect is less prominent for larger networks. One possibility is that the extra MLP block after the last Aggregation layer contributes far less than the rest of the network. Analyzing higher input sizes and higher image resolutions can shed more light on this. We have shown the best metrics in each column in bold-face, for both the cases with and without global features.
Fig. (3) shows the signal (top) efficiency vs. the background (QCD) rejection curves for those representative models for which at least one of the metrics is highest among others (those in bold-face in table 2).
In the case of other CNNs, the entries for the ‘Image Resolution’ column are calculated in the following way: we know these images were created $40\times 40$ (DeepTop) and $64\times 64$ (for ResNext) for $R=0.8$. This corresponds to $60\times 60$ and $96\times 96$ for $R=1.2$, which is our case. These are all grayscale images with only one channel for $p_{T}$. All networks used in this analysis that utilize the global features outperform the classic DeepTop CNN structure in terms of accuracy, if not in background rejection. ResNeXt-50 has a much better performance, but it is considerably heavier than the EffNet-S (the best performing global feat. variant is $1/7$-th in size of ResNeXt-50). These nets are faster as well. The longest inference time for EffNet-S is half of that for ResNeXt-50.
4 Conclusion
Jet tagging, a classification problem that aims to identify jets originating from different particles (top quarks, Higgs boson, gauge bosons, light quarks, or gluons), remains a challenging, albeit one of the most important tasks, especially in the context of a hadron collider like the LHC. Machine learning has now replaced traditional use of hand-crafted features to become an integral part of jet physics studies. Depending on the network used, jets are treated as images, sequences, trees or graphs. Existing literature shows that transformer based networks, which treat jets as particle clouds or graphs, perform the best. To achieve the highest possible accuracy, incorporating pairwise information is crucial. However, given the large number of pairwise interactions among the constituents of a jet, creating a fully connected graph becomes a computationally costly affair. CNNs, on the other hand, while being computationally low-cost, treat jets as images/arrays denoting the spatial arrangement of constituent particles in terms of energy, momentum, and mass; thus can utilize at least some local pairwise information. One can additionally incorporate global features of the jets to either directly to the images or to the network to enhance the jet identification accuracy. We have explored the possibility of combining these features and construct a computationally lightweight but efficient CNN-based network that can compete with other similar architectures.
With traditional CNNs, larger/deeper networks are needed to extract maximal information from the jet images, as opposed to typically lightweight CNNs like LeNet. Scaled up CNNs, such as ResNets are computationally costly. Using recent advancements in the construction of CNNs, we have used lightweight versions of EfficientNet to circumvent this issue. Although these were built to work on higher resolution images, one can follow a set of prescribed scaling rules to construct a network applicable for low resolution jet images as well. We have performed a comparative study with and without externally provided global features. Results for LeNet are also provided for comparison.
We observe that adding global features to the models, along with the images, results in better accuracy. Although this increase is somewhat monotonic with increasing image resolution for fully convolutional models, such as LeNet, the corresponding relationship for EffNet-S is not so straightforward. Increasing the input array size also improves the overall performance.
This exploration leads to three observations: first, there needs to be a systematic architecture search for versions of scalable models such as EfficientNet, tuned for jet classification; second, adding global jet features, in addition to the constituent information, makes it easier for smaller models to perform better; last, but not the least, as jets can be represented in multiple types of data, one can combine the strengths of different network architectures and build a hybrid ensemble of lightweight networks capable of achieving high accuracy while working with various representations of the same jet, e.g., image, constituent four-momenta, and global jet features. This ‘mixture of experts’ approach has the potential for superior performance, leading to better signal-to-background discrimination. Exploring this idea is beyond the scope of this paper, but it is something that we plan to work on as a subsequent project.
Acknowledgements.
RB, SM, and SKP would like to acknowledge support from ANRF, India (grant order no. CRG/2022/003208). SM would also like to acknowledge ANRF for grant order no.
CRG/2023/008570.

### v2

@@
-CNN on ‘Top’: In Search of Scalable & Lightweight Image-based Jet Taggers
+CNN on ‘Top’
 Abstract
@@
 We showcase the efficacy of our network in tagging top-quark jets in a sea of other light quark and gluon jets.
+The work also sheds light on the importance of global features for both the accuracy and the apparent redundancy of the network’s complexity.
 journal: Eur. Phys. J. C
@@
 As a result of this boost, the decay products of these standard model (SM) particles are highly collimated and can be contained within a single large-radius jet, which we call a ‘fat jet’.
-Scenarios like this can easily occur in various beyond the standard model (BSM) frameworks with vector-like quarks Yang:2025ktj ; Ghosh:2025gdq , leptoquarks Sahoo:2025kdj ; Bhaskar:2024snl ; Ghosh:2025gue , heavy right-handed neutrinos Chakraborty:2018khw ; Padhan:2022fak , supersymmetric particles with $R$-parity violating couplings Baruah:2024wrn ; Sahu:2025uop , to name a few.
+Scenarios like this can easily occur in various beyond the standard model (BSM) frameworks with vector-like quarks Yang et al.
+(2025); Ghosh et al.
+(2025a), leptoquarks Sahoo et al.
+(2025); Bhaskar and Mitra (2025); Ghosh et al.
+(2025b), heavy right-handed neutrinos Chakraborty et al.
+(2019); Padhan et al.
+(2022), supersymmetric particles with $R$-parity violating couplings Baruah et al.
+(2025); Sahu et al.
+(2025), to name a few.
 Among the above-mentioned SM particles, top quark holds a unique status.
@@
 Moreover, the top quark mass plays a crucial role in determining the stability of the electroweak vacuum.
-Though the current measurements suggest that our vacuum is ‘metastable’ Hiller:2024zjp ; Degrassi:2012ry ; Buttazzo:2013uya , this conclusion is highly sensitive to the precise values of the top quark measurements Alekhin:2012py ; Elias-Miro:2011sqh ; Espinosa:2015kwx .
+Though the current measurements suggest that our vacuum is ‘metastable’ Hiller et al.
+(2024); Degrassi et al.
+(2012); Buttazzo et al.
+(2013), this conclusion is highly sensitive to the precise values of the top quark measurements Alekhin et al.
+(2012); Elias-Miro et al.
+(2012); Espinosa (2016).
 Identifying top quark jets as fat jets in a sea of other light quark and gluon jets, therefore, is a rather important task.
-Machine learning (ML) has revolutionized this field Sahu:2024fzi ; Bhattacherjee:2022gjq ; Mikuni:2025tar by moving beyond traditional methods that relied on high-level, physics-inspired features like jet mass, jet radius, $n$-subjettiness, energy correlation functions, etc Thaler:2010tr ; Thaler:2011gf ; Stewart:2010tn ; Larkoski:2013eya ; Moult:2016cvt ; Komiske:2017aww ; Abdesselam:2010pt ; Kogler:2018hem ; Altheimer:2012mn ; Altheimer:2013yza ; Adams:2015hiv ; Marzani:2019hun ; Larkoski:2017jix .
+Machine learning (ML) has revolutionized this field Sahu et al.
+(2024); Bhattacherjee et al.
+(2024); Mikuni and Nachman (2025a) by moving beyond traditional methods that relied on high-level, physics-inspired features like jet mass, jet radius, $n$-subjettiness, energy correlation functions, etc Thaler and Van Tilburg (2011, 2012); Stewart et al.
+(2010); Larkoski et al.
+(2013); Moult et al.
+(2016); Komiske et al.
+(2018); Abdesselam and others (2011); Kogler and others (2019); Altheimer and others (2012, 2014); Adams and others (2015); Marzani et al.
+(2019); Larkoski et al.
+(2020).
 Deep learning models, on the other hand, instead of using simple, hand-crafted features, train themselves to represent the jets as points in a vector space of features best suited for the subsequent classification.
@@
 […]
-A more direct and fundamental approach to embedding is to represent the jet as a point cloud of its constituent particles.
+A more direct approach to embedding is to represent the jet as a point cloud of its constituent particles.
 Each ‘point’ is a particle with its own set of features, e.g., four momentum components, particle IDs, charges, and tracking information.
-GNNs Shlomi:2020gdn ; Duarte:2020ngm ; Thais:2022iok are particularly well-suited for this representation.
-The jet is now treated as a graph where each particle is a ‘node’, and the connections between them are ‘edges’ Komiske:2018cqr ; Qu:2019gqs ; Bogatskiy:2020tje ; Gong:2022lye ; Bogatskiy:2023nnw ; Bogatskiy:2022czk ; Bogatskiy:2023fug .
+GNNs Shlomi et al.
+(2020); Duarte and Vlimant (2020); Thais et al.
+(2022) are particularly well-suited for this representation.
+The jet is now treated as a graph where each particle is a ‘node’, and the connections between them are ‘edges’ Komiske et al.
+(2019); Qu and Gouskos (2020); Bogatskiy et al.
+(2020); Gong et al.
+(2022); Bogatskiy et al.
+(2024, 2022, 2023).
 Constituent information can also be represented as a varying length sequence, especially if it is ordered based on a certain property, e.g., by their transverse momentum $p_{T}$.
 RNNs—originally developed for processing text or speech—can then be used.
-The advent of the self- and cross-attention mechanisms gave us the famously parallelizable and scalable Transformer architecture DBLP:journals/corr/VaswaniSPUJGKP17 .
-Originally used to process and generate natural language sequences, these models can be thought of as special variations of GNNs JoshiTrans2GNN , and can process jets either as sequences of constituents or unordered point clouds.
+The advent of the self- and cross-attention mechanisms gave us the famously parallelizable and scalable Transformer architecture Vaswani et al.
+(2017).
+Originally used to process and generate natural language sequences, these models can be thought of as special variations of GNNs Joshi (2025), and can process jets either as sequences of constituents or unordered point clouds.
 One jarring issue that plagues the existing networks with the highest accuracy in identifying various jets is the computational cost.
@@
 It can either be gleaned by formidably large networks or can be preprocessed and fed to a smaller network later.
-The former is the path taken by transformer-based models Qu:2022mxj ; Mikuni:2021pou ; Wu:2024thh and more costly foundational models like Omnilearn(ed) Mikuni:2025tar ; Mikuni:2024qsr ; Bhimji:2025isp .
-Even these, when striving for higher accuracies, bring separately extracted features to the networks ‘attention’ [^1] Not adding pairwise information to the multi-head attention of the (Par-T) network results in a $1.2\%$ drop in accuracy and a $20-30\%$ drop in background rejection Qu:2019gqs ; Qu:2022mxj ..
+The former is the path taken by transformer-based models Qu et al.
+(2022); Mikuni and Canelli (2021); Wu et al.
+(2025) and more costly foundational models like Omnilearn(ed) Mikuni and Nachman (2025a, b); Bhimji et al.
+(2025).
+Even these, when striving for higher accuracies, bring separately extracted features to the networks ‘attention’ [^1] Not adding pairwise information to the multi-head attention of the (Par-T) network results in a $1.2\%$ drop in accuracy and a $20-30\%$ drop in background rejection Qu and Gouskos (2020); Qu et al.
+(2022)..
 In the latter approach, when jets are considered as particle clouds or graphs, creating a fully connected graph is prohibitively costly in terms of time and computation, mainly because the pairwise interactions (edge information222i.e., the abstract message-passing, not physical QCD interactions.) are numerous for a jet with a large number of constituents.
-These challenges are either bypassed by keeping only the nearest neighbor interactions (ParticleNet Qu:2019gqs ), or by representing data in a geometric algebra over space-time, equivariant under physics-motivated (Lorentz) transformations Bogatskiy:2022czk ; Bogatskiy:2023nnw ; Gong:2022lye .
-The newer architectures combine both of these approaches Spinner:2024hjm ; Brehmer:2024yqw .
+These challenges are either bypassed by keeping only the nearest neighbor interactions (ParticleNet Qu and Gouskos (2020)), or by representing data in a geometric algebra over space-time, equivariant under physics-motivated (Lorentz) transformations Bogatskiy et al.
+(2022, 2024); Gong et al.
+(2022).
+The newer architectures combine both of these approaches Spinner et al.
+(2024); Brehmer et al.
+(2025).
 Even running these state-of-the-art models on single-GPU machines becomes an ordeal, let alone training them from scratch.
@@
 Our objective in this work is to revisit computationally low-cost CNNs, as they implicitly utilize local, pairwise information (albeit spatially, through images) to some extent.
-Depending on the knowledge that the inclusion of global jet features enhances tagging efficiency Bhattacherjee:2022gjq , we explore the possibility of constructing a lightweight yet efficient network by incorporating global features with the images.
-Although computationally lightweight, smaller CNNs such as LeNet lenetPaper cannot achieve competitive accuracies.
+Depending on the knowledge that the inclusion of global jet features enhances tagging efficiency Bhattacherjee et al.
+(2024), we explore the possibility of constructing a lightweight yet efficient network by incorporating global features with the images.
+Although computationally lightweight, smaller CNNs such as LeNet Lecun et al.
+(1998) cannot achieve competitive accuracies.
 Extracting more information here is equivalent to increasing image resolution and/or model complexity (depth/width).
-Although CNNs become notoriously computationally costly when scaled up (e.g., ResNet ResNet1 ), recent developments, such as compound scaling EffNetV1 ; EffNetV2 , combined with depthwise and pointwise convolutionsMobileNetV1 ; MobileNetV2 ; MobileNetV3 ; MobileNetV4 ; MNASnet , resulted in networks that reach higher accuracies at a fraction of the cost.
+Although CNNs become notoriously computationally costly when scaled up (e.g., ResNet He et al.
+(2016)), recent developments, such as compound scaling Tan and Le (2019, 2021), combined with depthwise and pointwise convolutionsHoward et al.
+(2017); Sandler et al.
+(2018); Howard et al.
+(2019); Qin et al.
+(2024); Tan et al.
+(2019), resulted in networks that reach higher accuracies at a fraction of the cost.
 As all versions of EfficientNet were originally designed for images with higher resolution, we have built lightweight versions of those, while following the same scaling rules.
@@
 Finally, in Section 4, we summarize our work and draw conclusions.
-(a) $32\times 32$
-(b) $32\times 32$, Standardized
-(c) $32\xrightarrow{Crop}28$, Standardized
+(a) $35\times 35$
+(b) $35\times 35$, Standardized
+(c) $35\xrightarrow{Crop}28$, Standardized
 (d) $40\xrightarrow{Crop}32$, Standardized
-(e) $32\times 32$
-(f) $32\times 32$, Standardized
-(g) $32\xrightarrow{Crop}28$, Standardized
+(e) $35\times 35$
+(f) $35\times 35$, Standardized
+(g) $35\xrightarrow{Crop}28$, Standardized
 (h) $40\xrightarrow{Crop}32$, Standardized
 Figure 1: Representative images of two jets for different resolutions and pixel standardization.
-Although images in our dataset are rendered in RGB, we present them here using the CEI 1931 XYZ color scheme xyzcolorCEI:1931 , for better visual understanding.
+Although images in our dataset are rendered in RGB, we present them here using the CEI 1931 XYZ color scheme Smith and Guild (1931), for better visual understanding.
 All images from the top panel are of one representative top jet, and those below are of one QCD jet.
-The first column shows the jet constituents as a $32\times 32$ image in the $\Delta\eta-\Delta\phi$ plane, as the channels are combined and rendered.
+The first column shows the jet constituents as a $35\times 35$ image in the $\Delta\eta-\Delta\phi$ plane, as the channels are combined and rendered.
 The second panel in the same row contains the same image, with the pixels standardized (the mean image of the dataset is subtracted from the original and then divided by the standard deviation image).
@@
 The entire analysis is conducted in Mathematica® with the Wolfram Language, which uses the Apache MXNet® framework in the backend.
-The original EfficientNet architectures (before modification) were taken from the Wolfram Net Repository effnet:Wolfram .
+The original EfficientNet architectures (before modification) were taken from the Wolfram Net Repository M.
+Tan (2020).
 To compare the inference speed of CNNs in this work with some external models, inference times were calculated on another machine with an $11GB$ NVIDIA© GeForce GTX 1080-Ti graphics processor.
@@
 The objective, as mentioned before, is to identify top quark jets as signals and tag them as fat jets amid the sea of other jets.
-The events were generated with 14 TeV center-of-mass energy using Pythia8 Sjostrand:2014zea followed by detector simulation performed using Delphes deFavereau:2013fsa .
-The jets were constructed using the anti-kT algorithm Cacciari:2008gp with FastJet Cacciari:2011ma .
+The events were generated with 14 TeV center-of-mass energy using Pythia8 Sjöstrand et al.
+(2015) followed by detector simulation performed using Delphes de Favereau et al.
+(2014).
+The jets were constructed using the anti-kT algorithm Cacciari et al.
+(2008) with FastJet Cacciari et al.
+(2012).
 The pseudorapidity $|\eta_{j}|$ is required to be <2<2 for the fat top quark jet, whereas the radius parameter is chosen to be $0.8$.
@@
 To create each channel from the jet constituents, the hardest constituent is considered to be the center of the channel, and the rest of the image is constructed by binning the remaining constituents in two dimensions: the difference in pseudo-rapidity ($\Delta\eta$) and that in the azimuthal angle ($\Delta\phi$).
-The paper associated with the original dataset Kasieczka:2019dbj assumes the calorimeter resolution to be $0.04\times 2.25^{\circ}$, where a fat jet with a radius parameter $R=0.8$ can be covered with $40\times 40$ pixels.
-Fixing the bin sizes thus keeps our jet images at this dimension or less ($32\times 32$).
-Due to these choices and the nature of the dataset, both $\Delta\eta$ and $\Delta\phi$ axes are retained from 1.2-1.2 to $1.2$, without the loss of any constituents (see Figs (1(a)) and (1(e)) ).
-Unlike the original paper, we do not fix the axis of the second hardest object; rather, we use an image augmentation layer to randomly flip the image along either axis or both during training.
-We do not use random cropping, as the jets are constructed objects and can always be centered around the hardest $p_{T}$ constituent for all practical applications.
+The paper associated with the original dataset Butter and others (2019) assumes the calorimeter resolution to be $0.04\times 2.25^{\circ}$, where a fat jet with a radius parameter $R=0.8$ can be covered with $40\times 40$ pixels.
+Both $\Delta\eta$ and $\Delta\phi$ axes are retained from 1.2-1.2 to $1.2$ for this dataset, without the loss of any constituents.
+We keep more of the field of images than is usual practice in the community, where final images are generally cropped to $R=0.8$ Kasieczka et al.
+(2017); Butter and others (2019); Macaluso and Shih (2018); our images are center-cropped at $R=0.96$.
+To keep the images as small as possible, we use two sizes in the course of this work: $35\times 35$ (see Figs (1(a)) and (1(e)) ) and $40\times 40$.
+In the original paper, the axis of the image is rotated in such a way that the second hardest $p_{T}$ object falls on the positive $\Delta\phi$ axis and the third hardest constituent in the first quadrant.
+We have found out that this results in a lower accuracy for classification; rather, we use an image augmentation layer to randomly flip the image along either axis or both during training.
+We do not use random cropping, as the jets are constructed objects and can—for all practical applications—be centered around the hardest $p_{T}$ constituent.
 Furthermore, to use N-subjettiness later in our analysis, we also require each jet to have at least 4 constituents, which drops a meager 33 jets from the $201.8K$ dataset.
@@
 This clearly isolates a fairly circular region in the $\Delta\eta-\Delta\phi$ plane containing the bulk of the jet structure (see Figs (1(b)) and (1(f)) ).
-During some preliminary runs, it has been consistently observed that using images where the central region covers most of the image field yields better classification performance.
-Thus, we apply a center-crop to the $32\times 32$ images, making them $28\times 28$ (see Figs (1(c)) and (1(g)) ).
+Thus, we apply a center-crop to the $35\times 35$ images, making them $28\times 28$ ($R\sim 0.96$, see Figs (1(c)) and (1(g)) ).
 A similar size of the central region is obtained for the $40\times 40$ images when they are cropped to $32\times 32$ (see Fig.s (1(d)) and (1(h)) ).
-We keep more of the field of images than is usual practice in the community, where final images are generally cropped to $R=0.8$ Kasieczka:2017nvn ; Kasieczka:2019dbj ; Macaluso:2018tck .
-Our crops vary within $0.96\leq R\leq 1.05$.
 2.3 Data: Jet-Level Information
 We have combined some global features of the jet to ascertain how much they improve the tagging efficiency of the CNNs.
-The features used in this work are inspired by ref.s ATL-PHYS-PUB-2021-028 ; Bhattacherjee:2022gjq .
+The features used in this work are inspired by ref.s 38; B.
+Bhattacherjee, C.
+Bose, A.
+Chakraborty, and R.
+Sengupta (2024).
 The original dataset provides the four-momentum of up to 200 constituents per jet, from which all global features can be computed.
 For this purpose, we employ the fastjet package along with EnergyCorrelator contrib from fastjet-contrib-1.3.2.
-In addition to the global four-momentum of the jet, our feature set includes the N-subjettiness variables Thaler:2010tr , as well as several series of observables constructed from Energy Correlation Functions Larkoski:2013eya and generalized correlator functions Moult:2016cvt .
+In addition to the global four-momentum of the jet, our feature set includes the N-subjettiness variables Thaler and Van Tilburg (2011), as well as several series of observables constructed from Energy Correlation Functions Larkoski et al.
+(2013) and generalized correlator functions Moult et al.
+(2016).
 The complete list of variables is as follows:
@@
   • $C$-series: $C_{1}^{\beta}$, $C_{2}^{\beta}$, $C_{3}^{\beta}$ for $\beta=1.0,\,2.0$
-  • $D$-series [^4] The parameter choices for the D-series features follow Larkoski:2014gra ; Larkoski:2015kga ; Larkoski:2014zma .: $D_{2}^{\alpha,\beta}$ for $\alpha,\beta=1.0,\,2.0$, and the four $D_{3}$ variables defined in Bhattacherjee:2022gjq
+  • $D$-series [^4] The parameter choices for the D-series features follow Larkoski et al.
+(2014, 2016, 2015).: $D_{2}^{\alpha,\beta}$ for $\alpha,\beta=1.0,\,2.0$, and the four $D_{3}$ variables defined in Bhattacherjee et al.
+(2024)
   • U-series: $U_{1}^{\beta},U_{2}^{\beta},U_{3}^{\beta}$ for $\beta=0.5,\,1.0,\,2.0$
@@
   • N-series: $N_{2}^{\beta},N_{3}^{\beta}$ for $\beta=1.0,\,2.0$
-  • L-series: $L_{2},L_{3}$ as defined in ATL-PHYS-PUB-2021-028
+  • L-series: $L_{2},L_{3}$ as defined in 38
 Certain global features, such as $U_{3}^{\beta}$ and $M_{3}^{\beta}$, require at least 4 constituents for their computation.
 As mentioned in the previous sub-section, we discard all jets with $N_{\text{constituents}}<4$.
-For complete definitions and a detailed description of these variables, we refer the reader to ref.s Bhattacherjee:2022gjq ; ATL-PHYS-PUB-2021-028
+For complete definitions and a detailed description of these variables, we refer the reader to ref.s B.
+Bhattacherjee, C.
+Bose, A.
+Chakraborty, and R.
+Sengupta (2024); 38
 2.4 Model Architecture
-Historically, one of the most successful convolutional neural networks (CNNs), LeNet lenetPaper , is also the oldest and one of the simplest.
+Historically, one of the most successful convolutional neural networks (CNNs), LeNet Lecun et al.
+(1998), is also the oldest and one of the simplest.
 LeNet-5, with alternating convolutional and pooling layers, has been the dependable standard for smaller images (see the left side of Fig.
@@
 Over the decades following LeNet, CNNs continued to scale up to classify images at increasingly higher resolutions.
-AlexNet AlexNetPaper , Inception nets InceptionV1 ; InceptionV2V3 , ResNets ResNet1 and their wider versions wideResNet , DenseNets DenseNet , and MobileNets MobileNetV1 ; MobileNetV2 ; MobileNetV3 ; MobileNetV4 are only a few examples of these.
-To go deeper and wider, these newer architectures used varied techniques, such as ‘residual’ connections to feed information from earlier layers to the latter parts of the network, point-wise $(1\times 1)$ and depthwise convolutions for dimension reduction to remove computational bottlenecks, batch normalization, and mixtures of these methods InceptionV4 ; ResNext .
-Targeting general-purpose image classification, these models were optimized for the ImageNet dataset ImageNet2019 ; ImageNet2021 and its equivalents; thus, computationally, they are quite expensive.
+AlexNet Krizhevsky et al.
+(2012), Inception nets Szegedy et al.
+(2015, 2016), ResNets He et al.
+(2016) and their wider versions Zagoruyko and Komodakis (2016), DenseNets Huang et al.
+(2017), and MobileNets Howard et al.
+(2017); Sandler et al.
+(2018); Howard et al.
+(2019); Qin et al.
+(2024) are only a few examples of these.
+To go deeper and wider, these newer architectures used varied techniques, such as ‘residual’ connections to feed information from earlier layers to the latter parts of the network, point-wise $(1\times 1)$ and depthwise convolutions for dimension reduction to remove computational bottlenecks, batch normalization, and mixtures of these methods Szegedy et al.
+(2017); Xie et al.
+(2017).
+Targeting general-purpose image classification, these models were optimized for the ImageNet dataset Yang et al.
+(2020, 2022) and its equivalents; thus, computationally, they are quite expensive.
 Layers Input Resolution
@@
 Table 1: Effnet-S configurations for two different input resolutions. For more details, see sec. 2.4.
-Network Input Image No. of Inference Time Accuracy AUC Bkg. Rej at TPR
-Type Array Size Resolution Param.s /100 Jets ($ms$) (%) (%) $50\%$ $30\%$
-LeNet $3\times 28\times 28$ $32\to 28$ $428K$ $5.5(7)$ $92.52$ $97.81$ $160(2)$ $525(23)$
-(Images) $3\times 32\times 32$ $653K$ $7.3(8)$ $92.54$ $97.80$ $162(2)$ $536(14)$
-$3\times 40\times 40$ $1253K$ $6.5(20)$ $92.56$ $97.82$ $169(3)$ $568(31)$
-$3\times 32\times 32$ $40\to 32$ $653K$ $6.7(6)$ $92.55$ $97.82$ $165(4)$ $516(20)$
-$3\times 40\times 40$ $1253K$ $6.5(7)$ $92.51$ $97.75$ $155(2)$ $476(16)$
-$3\times 28\times 28$ $32\to 28$ $485K$ $6.4(6)$ $93.12$ $98.11$ $225(4)$ $774(33)$
-$3\times 32\times 32$ $710K$ $7.5(5)$ $93.10$ $98.11$ $222(6)$ $777(35)$
-(Images + $3\times 40\times 40$ $1310K$ $7.3(7)$ $93.08$ $98.11$ $222(3)$ $795(23)$
-Global Feat.s) $3\times 32\times 32$ $40\to 32$ $710K$ $7.0(7)$ $93.17$ $98.14$ $233(2)$ $793(19)$
-$3\times 40\times 40$ $1310K$ $7.7(8)$ $\mathbf{93.20}$ $\mathbf{98.15}$ $237(3)$ $\mathbf{814(24)}$
-EffNet-S $3\times 40\times 40$ $32\to 28$ $162K$ $7.7(7)$ $92.59$ $97.85$ $174(5)$ $\mathbf{602(14)}$
-(Images) $3\times 64\times 64$ $305K$ $10.4(6)$ $\mathbf{92.70}$ $\mathbf{97.91}$ $170(13)$ $577(67)$
-$3\times 40\times 40$ $40\to 32$ $162K$ $8.5(6)$ $91.97$ $97.51$ $129(9)$ $415(40)$
-(Images + $3\times 40\times 40$ $32\to 28$ $208K$ $8.4(6)$ $93.12$ $98.13$ $233(6)$ $\mathbf{864(44)}$
-Global Feat.s) $3\times 64\times 64$ $365K$ $12.0(11)$ $\mathbf{93.13}$ $98.15$ $226(8)$ $817(38)$
-DeepTop Macaluso:2018tck ; Kasieczka:2017nvn ; Kasieczka:2019dbj $1\times 37\times 37$ $60\to 40\to 37$ $610K$ - $93.00$ $98.10$ - $914(14)$
-ResNeXt-50 ResNext
-Ref. Kasieczka:2019dbj $1\times 64\times 64$ $96\to 64$ $1460K$ - $93.60$ $98.40$ - $1122(47)$
-Ref. Qu:2019gqs $1\times 64\times 64$ $96\to 64$ $1460K$ 22 $93.60$ $98.37$ $302(5)$ $1147(58)$
+Network Input Img. Res. No. of Inference Accuracy AUC Bkg. Rejection
+Type Array ($Org\to$ Param.s Time ($ms$) (%) at TPR
+($3\times n\times n$) CrpCrp) Param.s /100 Jets Avg. Best Avg. Best $50\%$ $30\%$
+$28$ $35\to 28$ $428K$ $5.7(8)$ $92.42(3)$ $92.48$ $97.74(1)$ $97.76$ $156(2)$ $480(18)$
+LeNet $32$ $653K$ $6.1(9)$ $92.40(2)$ $92.43$ $97.74(1)$ $97.74$ $154(3)$ $483(13)$
+(Images) $40$ $1253K$ $6.4(8)$ $92.45(3)$ $\mathbf{92.50}$ $97.76(2)$ $\mathbf{97.78}$ $\mathbf{161(4)}$ $\mathbf{514(23)}$
+$32$ $40\to 32$ $653K$ $11.9(10)$ $92.41(4)$ $92.48$ $97.75(1)$ $97.76$ $161(3)$ $512(25)$
+$40$ $1253K$ $13.4(18)$ $92.42(5)$ $92.48$ $97.74(1)$ $97.74$ $154(3)$ $476(11)$
+$28$ $35\to 28$ $485K$ $5.7(8)$ $93.33(2)$ $93.37$ $98.24(0)$ $98.24$ $258(5)$ $916(25)$
+(Images $32$ $710K$ $6.1(8)$ $93.34(2)$ $\mathbf{93.38}$ $98.24(1)$ $\mathbf{98.26}$ $257(3)$ $942(31)$
++ $40$ $1310K$ $6.6(9)$ $93.31(3)$ $93.36$ $98.23(1)$ $98.25$ $255(5)$ $909(34)$
+Global $32$ $40\to 32$ $710K$ $14.5(16)$ $93.33(2)$ $93.36$ $98.24(0)$ $98.24$ $255(2)$ $\mathbf{951(43)}$
+Feat.s) $40$ $1310K$ $14.8(7)$ $93.34(2)$ $93.37$ $98.24(1)$ $98.25$ $\mathbf{263(3)}$ $906(32)$
+$36$ (S0) $35\to 28$ $158K$ $7.3(9)$ $92.31(11)$ $92.50$ $97.73(4)$ $97.80$ $160(5)$ $536(24)$
+EffNet-S $40$ (S1) $184K$ $8.0(7)$ $92.35(8)$ $92.46$ $97.75(3)$ $97.76$ $163(3)$ $552(18)$
+$48$ (S2) $215K$ $8.7(8)$ $92.39(9)$ $92.53$ $97.76(4)$ $97.80$ $\mathbf{167(5)}$ $569(30)$
+$56$ (S3) $258K$ $9.6(8)$ $92.58(1)$ $92.72$ $97.85(3)$ $97.90$ $171(6)$ $\mathbf{585(32)}$
+(Images) $64$ (S4) $305K$ $10.6(8)$ $92.20(40)$ $\mathbf{92.75}$ $97.74(11)$ $\mathbf{97.92}$ $156(16)$ $499(78)$
+$36$ (S0) $40\to 32$ $158K$ $13.5(10)$ $91.40(60)$ $91.91$ $97.37(13)$ $97.50$ $127(11)$ $431(49)$
+$40$ (S1) $184K$ $13.9(10)$ $90.5(12)$ $91.80$ $97.22(26)$ $97.39$ $112(15)$ $361(57)$
+$48$ (S2) $215K$ $14.8(10)$ $91.4(10)$ $92.05$ $97.35(23)$ $97.56$ $126(18)$ $407(64)$
+$56$ (S3) $258K$ $15.8(7)$ $92.0(4)$ $92.31$ $97.60(10)$ $97.68$ $143(9)$ $460(45)$
+$64$ (S4) $305K$ $16.5(9)$ $91.9(6)$ $92.42$ $97.52(18)$ $97.74$ $137(17)$ $434(63)$
+$36$ (S0) $35\to 28$ $242K$ $8.2(9)$ $93.28(2)$ $93.32$ $98.21(1)$ $\mathbf{98.23}$ $242(5)$ $890(34)$
+(Img.s + $40$ (S1) $276K$ $7.9(9)$ $93.26(2)$ $93.30$ $98.21(1)$ $98.22$ $241(7)$ $859(34)$
+Glob. Feat.s) $48$ (S2) $315K$ $8.4(9)$ $93.20(2)$ $93.30$ $98.18(2)$ $98.20$ $241(10)$ $839(51)$
+$56$ (S3) $368K$ $10.0(10)$ $93.31(5)$ $\mathbf{93.38}$ $98.21(1)$ $98.22$ $232(5)$ $817(26)$
+$64$ (S4) $422K$ $11.4(9)$ $93.15(5)$ $93.24$ $98.11(3)$ $98.15$ $208(8)$ $707(48)$
+$36$ (S0) $40\to 32$ $242K$ $13.8(8)$ $93.24(8)$ $93.31$ $98.20(4)$ $\mathbf{98.23}$ $244(13)$ $\mathbf{897(74)}$
+$40$ (S1) $276K$ $13.6(12)$ $93.26(2)$ $93.31$ $98.20(1)$ $98.21$ $\mathbf{247(6)}$ $868(25)$
+$48$ (S2) $315K$ $14.4(9)$ $93.20(3)$ $93.26$ $98.17(2)$ $98.20$ $235(6)$ $864(30)$
+$56$ (S3) $368K$ $15.7(9)$ $93.22(9)$ $93.35$ $98.18(3)$ $98.23$ $237(9)$ $829(44)$
+$64$ (S4) $422K$ $17.0(10)$ $93.21(5)$ $93.27$ $98.15(3)$ $98.18$ $225(13)$ $778(71)$
+DeepTop $1\times$ $60\to 40$
+Macaluso and Shih (2018); Kasieczka et al. (2017); Butter and others (2019) $\times 37 37\to 37$ $610K$ $93.00$ $98.10$ $914(14)$
+ResNeXt-50 Xie et al. (2017)
+Ref. Butter and others (2019) $1\times$
+$64\times 64$ $96\to 64$ $1460K$ $93.60$ $98.40$ $1122(47)$
+Ref. Qu and Gouskos (2020) $1\times$
+$64\times 64$ $96\to 64$ $1460K$ 22 $93.60$ $98.37$ $302(5)$ $1147(58)$
 Table 2: Single number performance metrics for ML models performed on the total test dataset.
 From the second to the seventh column, we respectively report the input array sizes, the resolution of the generated jet images ($A\to B$ means the original image was $A\times A$, and it was cropped to $B\times B$), total number of parameters in the network, inference time for 100 jets in milliseconds, overall accuracy, and area under the ROC curve (AUC).
-The last two columns list the QCD background rejection at top tagging efficiencies of $50\%$ and $30\%$ respectively.
-Though scaling CNNs according to image resolution has been attempted before long2015fully , EfficientNets EffNetV1 ; EffNetV2 were the first to solve the scaling problem by using ‘compound scaling’, connecting network depth and width to image resolution.
+The last two columns list the QCD background rejection at top tagging efficiencies of $50\%$ and $30\%$, respectively.
+The values in bold-face represent the best metrics for a specific type of Network.
+Though scaling CNNs according to image resolution has been attempted before Long et al.
+(2015), EfficientNets Tan and Le (2019, 2021) were the first to solve the scaling problem by using ‘compound scaling’, connecting network depth and width to image resolution.
 Given a baseline network (found from a neural architecture search, or NAS), these quantities scale as:
@@
 $\text{resulution coefficient }r$ $=\gamma^{\phi}\,$, (1)
+(a) Model Complexity vs Best Accuracy
+(b) Model Complexity vs Accuracy Dist.
+(c) Plot Legend
+(d) Size Comparison between EffNet-B, EffNet-S, and LeNet
+Figure 3: Performance comparison of different CNN architectures: fig.
+(a) shows the best accuracies vs.
+number of model parameters of the nets used, fig.
+(b) shows the median accuracies and error-bars in the same setting, fig.
+(c) is the legend for (a) and (b), and fig.
+(d) is a comparison of sizes of the networks compared to traditional EfficientNet versions.
 where $\alpha\geq 1$, $\beta\geq 1$, $\gamma\geq 1$.
@@
 There are seven main building blocks of an EfficientNet, each containing multiple parts.
-Building upon other successful lightweight models MobileNetV2 ; MNASnet , EfficientNets replace a fully convolutional operation with a depthwise convolution (a single filter per channel) and a pointwise ($1\times 1$) convolution (see the blown up schematic of Block 2A in Fig.
+Building upon other successful lightweight models Sandler et al.
+(2018); Tan et al.
+(2019), EfficientNets replace a fully convolutional operation with a depthwise convolution (a single filter per channel) and a pointwise ($1\times 1$) convolution (see the blown up schematic of Block 2A in Fig.
 (2) ).
 This significantly reduces the computational cost while maintaining similar performance.
-When stacked together, the blocks mentioned above also perform a function known as inverted bottlenecking, embedding the input manifold into a lower dimensional manifold.
-Each of these blocks also contains a squeeze and expansion block squeeze , with a residual connection maintaining the information flow, as well as a combination of Swish, RELU, and Sigmoid activation functions.
-More details can be found in the original paper EffNetV1 .
+When stacked together, the blocks mentioned above also perform a function known as inverted bottlenecking, embedding the input manifold into a lower-dimensional manifold.
+Each of these blocks also contains a squeeze and expansion block Hu et al.
+(2018), with a residual connection maintaining the information flow, as well as a combination of Swish, RELU, and Sigmoid activation functions.
+More details can be found in the original paper Tan and Le (2019).
 While the width coefficient scales up the channel dimension of each block, the depth coefficient determines the number of times each type of block is repeated, making the network deeper.
@@
 The solid lines are for models working on both images and global features, dashed ones correspond to those with only images as input.
+The solid black curve is the optimal ROC curve obtainable by any model, obtained from likelihood ratio analysis (ref.
+Geuskens:2024tfo).
 To circumvent this without changing the hyperparameters of the original EfficientNet architecture (and using the prescribed values of $\alpha$, $\beta$, and $\gamma$), we use negative $\phi$ values to scale down the input resolution.
@@
 Any smaller, and the bottlenecks within the network would not hold without modification.
-We choose the starting resolution to be $40\times 40$ (in accordance with the resolution of the simulated calorimeter), corresponding to $\phi\simeq-12.33$.
-A larger version was also created with $64\times 64$ inputs, for $\phi\simeq-8.96$ [^6] Similar efforts can be found here: https://github.com/qubvel/efficientnet/issues/26.
+We choose the starting resolution to be $36\times 36$, corresponding to $\phi=-13$, then repeatedly increase $\phi$ by 1 to get the rest of the configurations.
+We thus create 5 variants, the last one corresponding to $\phi=-9$ [^6] Similar efforts can be found here: https://github.com/qubvel/efficientnet/issues/26.
 The other important parameter to set is the dropout probability, which is gradually increased from $B_{0}$ to $B_{7}$, as more regularization is needed for larger networks.
-We linearly extrapolated these values using the width and depth coefficients and used the corresponding values for our networks.
-Evidently, the much smaller size of our networks resulted in significantly less dropout regularization.
+We linearly extrapolated these values using the width and depth coefficients and found that for input array widths <90<90, we do not need dropouts.
 Altogether, we will denote this smaller architecture as EfficientNet-Small (EffNet-S) for this work.
@@
 An image of the trained network is saved from each round and used in the next one.
-This is done for 15 (18) rounds for $32\times 32$ ($40\times 40$) image representation of jets to ensure that the entire dataset is adequately represented.
+This is done for 12 rounds for $35\times 35$ ($40\times 40$) image representation of jets to ensure that the entire dataset is adequately represented.
 For most of the models used in this work, we repeat the process, starting with the last trained network, and use the same rounds again.
-This way, the rounds mimic the behavior of epochs, while storing the datasets beforehand reduces computation time.
-We then evaluate the accuracy of each of these models on the test dataset, and the one with the highest accuracy is chosen as the best version of that specific model architecture.
+This way, the rounds mimic the behavior of epochs, while storing the datasets in advance reduces computation time.
+We then evaluate the accuracy of each model on the test dataset, and the model with the highest accuracy is saved as the best version of that specific architecture.
 All further measurements are performed on this iteration of the model.
 To ensure optimal computation time, we choose batch sizes depending on whether the memory dedicated to the graphics card can handle them without overflowing into the SWAP memory.
-For EffNet-S (64) a batch size of $2048$ is used.
+For EffNet-S versions 0, 11, and $2$, a batch size of $4096$ is used; for EffNet-S versions $3$ and $4$, these numbers are respectively $3072$ and $2560$.
 For all other networks, it is fixed at $4096$.
-We use the ADAM optimizer kingma2014adam with default settings ($\beta_{1}=0.9$, $\beta_{2}=0.99$, and $\epsilon=10^{-5}$, along with the default learning rate update schedule) to minimize the cross-entropy loss, where we wait for a fixed number of epochs (generally 10) for the last minimum loss to be updated; otherwise, we stop the training.
+We use the ADAM optimizer Kingma (2014) with default settings ($\beta_{1}=0.9$, $\beta_{2}=0.99$, and $\epsilon=10^{-5}$, along with the default learning rate update schedule) to minimize the cross-entropy loss, where we wait for a fixed number of epochs (generally 10) for the last minimum loss to be updated; otherwise, we stop the training.
 The global features of the jets are standardized and utilized in the second part of the analysis, in conjunction with the images.
@@
 Keeping that in mind, we use the first part of the best network of a particular architecture and add an MLP block to it in this part of the analysis.
-We retrain the best network of each architecture nine additional times, with a representative subset of the dataset, to obtain the uncertainties arising from random weight initialization.
+We retrain the best network of each architecture 9 additional times, with a representative subset of the dataset, to obtain the uncertainties arising from random weight initialization.
 3 Results
@@
 The table is divided into three blocks: the first block contains all the information about the LeNet architecture, the second one about EffNet-S, and the third one lists some of the publicly available information about CNNs used previously as top-taggers for comparison.
-We have obtained results for LeNet-type networks with three different input resolutions: $28\times 28$, $32\times 32$, and $40\times 40$.
-Versions of EffNet-S are of two types: $40\times 40$ and $64\times 64$.
-Images are generated in two ways: $32\times 32$ images cropped to $28\times 28$ and $40\times 40$ images cropped to $32\times 32$, as explained in sec.
+We have obtained results for LeNet-type networks with three two different input resolutions: $35\times 35$, and $40\times 40$.
+Versions of EffNet-S are of 5 types: 0 to $4$.
+Images are generated in two ways: $35\times 35$ images cropped to $28\times 28$ and $40\times 40$ images cropped to $32\times 32$, as explained in sec.
 2.2.
@@
 The LeNet-40 is $2.9(2.7)$ times the size of LeNet-40 without(with) global features.
-The smallest EffNet-S, while applied only to images, outperforms all LeNets, while being $1/8$-th the size of the largest LeNet and in terms of the total number of parameters.
-Its performance is best when the input images are $32\xrightarrow{Crop}28$, and it drops a little for the larger images.
+The smallest EffNet-S, while applied only to images, achieves accuracy similar to the largest LeNet, while being $1/8$-th in size (the total number of parameters).
+Its performance is best when the input images are $35\xrightarrow{Crop}28$, but it drops quite a lot for the larger images.
+Adding global features is beneficial here as well, but the effect is less prominent for larger networks.
+We have shown the best metrics in each column in bold-face, for both the cases with and without global features.
+As was mentioned earlier in sec.
+2.1, the inference timings were calculated in a separate machine with a weaker GPU.
+While networks consistently take more time for larger image sizes, all of the networks dealt with in this work have almost half the inference time of that of ResNeXt-50.
+Fig.
+(3) summarizes the accuracies of table (2) in comparison to architecture complexity (number of net-parameters).
+Fig.
+(3(d)) shows the relative sizes of different versions of EffNet-S compared to the versions of LeNets used in this analysis as well as the larger and standard EfficientNets (B0 to B7), where the number of parameters is in log-scale.
+Even the largest EffNet-S is smaller than the smallest LeNet and all of these are an order of magnitude smaller than EffNet-Bs.
+In figures (3(a)) and (3(b)), accuracies of the nets are compared, with their model complexity in the $x$-axis.
+While fig.
+(3(a)) plots the best accuracies for each type of net, fig.
+(3(b)) shows the variation in it arising from random weight initialization.
+These employ 10 runs each, as mentioned at the end of the last section.
+As far as accuracies are concerned and for only images as input, we see that the performance of LeNet does not change much (and remains within randomization error), with increasing complexity, for all image resolutions.
+The EffNet-S, on the other hand, performs worse for larger image sizes.
+For the larger variants, this loss in performance falls within the randomization error-bar, but overall, the performance not only worsens, but has larger variance for $40\xrightarrow{Crop}32$ images.
 The possible reason is that the ‘MB-Conv’ blocks with bottlenecks in EffNet-S do not perform at the level of full convolutions when analyzing sparse pixel-level information, and networks with larger input arrays are needed for proper classification.
-Adding global features is beneficial here as well, but the effect is less prominent for larger networks.
-One possibility is that the extra MLP block after the last Aggregation layer contributes far less than the rest of the network.
+While the addition of global features improves network performances by a margin, the accuracies become more or less independent of the network complexity across architectures, especially when randomization errors are considered.
+One possibility is that the extra MLP block after the last Aggregation layer contributes far less than the rest of the network, but the more probable explanation would be that the impact of global features overshadow the CNN-performance.
 Analyzing higher input sizes and higher image resolutions can shed more light on this.
-We have shown the best metrics in each column in bold-face, for both the cases with and without global features.
+More work is needed regarding the integration of the global features in these networks.
 Fig.
@@
 the background (QCD) rejection curves for those representative models for which at least one of the metrics is highest among others (those in bold-face in table 2).
+This also shows how all EffNet-S nets outperform the best LeNet with just images as input (dashed curves), but with Global features put in, all of their performances are equivalent (solid, colored curves).
+We have also added the optimal possible ROC curve obtained from likelihood ratio analysis using generative models (ref.
+Geuskens:2024tfo; in thick black) to showcase the space allowed for future progress.
 In the case of other CNNs, the entries for the ‘Image Resolution’ column are calculated in the following way: we know these images were created $40\times 40$ (DeepTop) and $64\times 64$ (for ResNext) for $R=0.8$.
@@
 Depending on the network used, jets are treated as images, sequences, trees or graphs.
-Existing literature shows that transformer based networks, which treat jets as particle clouds or graphs, perform the best.
+Existing literature shows that transformer-based networks, which treat jets as particle clouds or graphs, perform the best.
 To achieve the highest possible accuracy, incorporating pairwise information is crucial.
@@
 With traditional CNNs, larger/deeper networks are needed to extract maximal information from the jet images, as opposed to typically lightweight CNNs like LeNet.
-Scaled up CNNs, such as ResNets are computationally costly.
+Scaled-up CNNs, such as ResNets are computationally costly.
 Using recent advancements in the construction of CNNs, we have used lightweight versions of EfficientNet to circumvent this issue.
@@
 We have performed a comparative study with and without externally provided global features.
-Results for LeNet are also provided for comparison.
+Results for LeNet are also provided as a benchmark for comparison.
 We observe that adding global features to the models, along with the images, results in better accuracy.
-Although this increase is somewhat monotonic with increasing image resolution for fully convolutional models, such as LeNet, the corresponding relationship for EffNet-S is not so straightforward.
-Increasing the input array size also improves the overall performance.
-This exploration leads to three observations: first, there needs to be a systematic architecture search for versions of scalable models such as EfficientNet, tuned for jet classification; second, adding global jet features, in addition to the constituent information, makes it easier for smaller models to perform better; last, but not the least, as jets can be represented in multiple types of data, one can combine the strengths of different network architectures and build a hybrid ensemble of lightweight networks capable of achieving high accuracy while working with various representations of the same jet, e.g., image, constituent four-momenta, and global jet features.
-This ‘mixture of experts’ approach has the potential for superior performance, leading to better signal-to-background discrimination.
-Exploring this idea is beyond the scope of this paper, but it is something that we plan to work on as a subsequent project.
+Although this the increase in performance is somewhat monotonic with increasing image resolution for fully convolutional models, such as LeNet, the corresponding relationship for EffNet-S is not so straightforward; increasing the input array size also improves the overall performance.
+The performance of these models becomes somewhat independent of their architecture in the presence of the global features.
+When applied to jet-classification tasks, this begs for further inquiry into the interplay between global features and constituent-level information.
+This exploration leads to three observations: first, there needs to be a systematic architecture search for versions of scalable models such as EfficientNet, tuned for jet classification; second, adding global jet features, in addition to the constituent information, makes it easier for smaller models to perform better, but the exact way these features are fed-in needs further exploration; last but not least, as jets can have multiple types of representation, e.g., image, constituent four-momenta, and global jet features, one can combine the strengths of different network architectures and build a hybrid ‘mixture of experts’ ensemble of lightweight networks with the potential for superior performance, leading to better signal-to-background discrimination.
+Exploring these ideas is beyond the scope of this paper, but it is something that we plan to work on as a subsequent project.
 Acknowledgements.